OpenStudy (anonymous):

solve using the square root method. Round the solution to the nearest tenth. x^2-20=0

OpenStudy (anonymous):

-4.5,4.5?

sam (.sam.):

x^2-20=0 x^2-20+20=20 x^2=20 \[\sqrt{x^2}=\sqrt{20}\] \[x=\sqrt{20}\] \[x=\sqrt{4}\sqrt{5}\] \[x=2\sqrt{5}\]

sam (.sam.):

yes there should be \[\huge \pm \] when square rooting, So its -4.5,4.5